OpenStudy (anonymous):

2. How is the graph of y = –2x² – 5 different from the graph of y = –2x²? (1 point) \It is A.shifted 5 units up. B.It is shifted 5 units down. C.It is shifted 5 units left. D.It is shifted 5 units right.

OpenStudy (anonymous):

It's b

OpenStudy (anonymous):

y = –2x² – 5 Put 0 in for x y = -2(0)^2 - 5 y = -5 y = –2x² put 0 in for x y = -2(0) y = 0